OpenStudy (anonymous):

How do you show y=x-10

OpenStudy (anonymous):

y=x-10 x=y+10 y=x-10

OpenStudy (anonymous):

what do you mean

OpenStudy (anonymous):

I don't know what your problem wants me to do, sorry.

OpenStudy (anonymous):

did you mean show by graphing?